OpenStudy (anonymous):

Find the slope of the line connecting these points. (4, 2), (8, 0)

OpenStudy (aroub):

0-2/8-4 -2/4 -1/2

OpenStudy (anonymous):

Find the slope of the line connecting these points. (2, 0), (10, 2)

OpenStudy (aroub):

2-0/10-2 2/8 1/4

OpenStudy (anonymous):

18 medal

OpenStudy (anonymous):

Find the slope of the line connecting these points. (8, 10), (4, 3)

OpenStudy (aroub):

3-10/4-8 -7/-4 7/4

OpenStudy (anonymous):

Find the slope of the line connecting these points. (6, 4), (4, -3)

OpenStudy (aroub):

wait wait from where are u getting these questions??

OpenStudy (anonymous):

my summer school submissions

OpenStudy (aroub):

lol oh okay!! wuld u like to try to solve one ??

OpenStudy (anonymous):

i dont have time ive been trying

OpenStudy (aroub):

okay!

OpenStudy (aroub):

-3-4/4-6 -7/-2 7/2

OpenStudy (anonymous):

Find the slope of the line connecting these points. (7, 5), (4, 9)

OpenStudy (aroub):

9-5/4-7 4/-3

OpenStudy (anonymous):

Find the slope of the line connecting these points. (-5, 9), (7, -6)

OpenStudy (aroub):

-6-9/7+5 -15/12 -5/4

OpenStudy (anonymous):

Find the slope of the line connecting these points. (-8, -7), (6, -3)

OpenStudy (aroub):

-3+7/6+8 4/14 2/7

OpenStudy (anonymous):

Find the slope of the line connecting these points. (-9, 0), (3, 5)

OpenStudy (aroub):

oh my how many questions are they??

OpenStudy (anonymous):

theres like 4 more

OpenStudy (aroub):

okay

OpenStudy (aroub):

5-0/3+9 5/12

OpenStudy (anonymous):

Find the x- and y- intercepts for the graph of the following equation. 7x + 6y = 126

OpenStudy (aroub):

ok, if u want to find the x-intercept u put y=0 and when u want to find the y-intercept u put x=0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

x-intercept is 18, y-intercept is 21

OpenStudy (anonymous):

Find the x- and y- intercepts for the graph of the following equation. 5x + 3y = -30

OpenStudy (anonymous):

just this one and then one men im doneore th

OpenStudy (anonymous):

i ment just that one and then 1 more then im done

OpenStudy (anonymous):

x-intercept = -6,y-intercepts = -10

OpenStudy (aroub):

okay x-inter is -6 and y-inter is -10

OpenStudy (anonymous):

Find the x- and y- intercepts for the graph of the following equation. 3x + 7y = -63

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

last one brahhhh chilll

OpenStudy (aroub):

x-inter is -21 and y-inter is -9

OpenStudy (aroub):

done?!?!

OpenStudy (anonymous):

yessss thanks u

OpenStudy (anonymous):

medal

OpenStudy (aroub):

oh yea , ANYTIME! :D